# Meta-Learning Driven Movable-Antenna-assisted Full-Duplex RSMA for Multi-User Communication: Performance and Optimization

### v2

@@ Abstract
Full-duplex (FD) radios at base station (BS) have gained significant interest because of their ability to simultaneously transmit and receive signals on the same frequency band. However, FD communication is hindered by self-interference (SI) and intra-cell interference caused by simultaneous uplink (UL) transmissions affecting downlink (DL) reception. Movable antenna (MA) technology has recently emerged as a promising solution to mitigate interference by dynamically adjusting antenna positions within the transmitter or receiver region, enhancing multiplexing gain and spectral efficiency (SE). In traditional FD systems with fixed-position antennas (FPAs), residual SI and intra-cell interference are mitigated either through beamforming and power allocation or by leveraging the inherent interference management strategy of multiple access schemes like rate-splitting multiple access (RSMA). Utilizing MA in FD systems adds an extra degree of freedom to effectively mitigate interference in conjunction with other strategies. how to write it nicely. In this context, we integrate MA with FD systems to evaluate and investigate its effectiveness in handling SI and intra-cell interference. Moreover, we utilize RSMA as our multiple access technique in both UL and DL transmission. To achieve the full potential of the system, we evaluate three different scenarios with FD-BS-RSMA with MAs where our goal is to maximize the total sum rate of the system by jointly optimizing the transmitting and receiving beamforming vectors, UL user equipment (UE) transmission power, MA positions, and common stream split ratio of RSMA while satisfying the minimum data rate requirements of all UEs, common stream rate constraint, power budget requirements of BS and UL UEs, and inter-MA distance. The formulated optimization problem is highly non-convex in nature, and hence, we propose a gradient-based meta-learning (GML) approach which can handle the non-convexity in a discrete manner by optimizing each variable in a different neural network. Our simulation results show that introducing MA with FD-BS-RSMA system not only improves the overall SE but also effectively tackle SI and intra-cell interference. Our proposed three scenarios, which integrate MA with the FD-BS-RSMA system, achieve performance improvements ranging from 41% to 240% compared to traditional FPA systems.
Index Terms:
Movable antenna, RSMA, full-duplex (FD), uplink, downlink, meta-learning
@@ I Introduction
With the emergence of sixth-generation (6G) mobile communication systems, there is a pressing demand to explore communication technologies that offer high spectral efficiency (SE), reliability, energy efficiency and latency [1]. In addition, the continuous surge in data-intensive applications places highly demanding requirements on wireless communication systems. However, due to limited spectrum resources, enhancing the communication rate using traditional half-duplex (HD) systems remains a challenge. In order to fulfill these stringent requirements and challenges, several promising technologies have emerged, such as millimeter wave (mmWave) communication, massive multiple-input multiple output (mMIMO) and full-duplex (FD) communications [1]. Among these technologies, FD communication has garnered significant interest because of its superior SE and co-time co-frequency transceiving capability compared to its HD counterpart [2].
@@
In FD communications, multiple terminals can simultaneously transmit and receive information within the same frequency band [3]. In particular, FD communication is capable of doubling up the SE by accommodating both uplink (UL) and downlink (DL) communications at the same time-frequency resource. However, allocating transmit and receive signals to operate in the same frequency and time slot inherently leads to significant interference issues. Specifically, the most critical issue is residual self-interference (SI) from the transmit antennas of the base station (BS) to the receive antennas, which could be much higher than the received signal. In addition, residual SI and DL user equipment (UE) experience significant intra-cell interference caused by UL UE transmissions, which greatly limits the potential gains of using FD technology. Hence, FD technology will greatly suffer quite a loss if the interferences are not handled in an efficient manner.
@@
Recently, movable antenna (MA) has emerged as a contender technology that is capable of effectively reducing interference in the MIMO system by positioning each MA in the transmitter/receiver region to shift from areas of high interference to locations with minimal interference, resulting in higher multiplexing gain and higher SE [4]. Moreover, by moving the antenna positions in a certain region, MA adds a new degree of freedom to exploit the benefit of the spatial domain. Specifically, unlike traditional fixed-position antennas (FPAs), MAs can be dynamically adjusted within a spatial region to optimize channel conditions and enhance communication performance. In particular, an MA is linked to the radio frequency (RF) chain via a flexible cable, enabling it to be repositioned within a spatial region with the help of a driver. Although traditional MIMO systems using antenna selection can achieve spatial degrees of freedom, they generally require a large number of antennas to cover the entire spatial region and fully leverage spatial diversity and multiplexing gains [4]. In contrast, a much smaller number of MAs, even a single MA, can fully exploit these degrees of freedom because of their flexible positioning.
@@
Another crucial aspect of enhancing SE lies in the careful selection of multiple access techniques. In the past couple of years, Rate-Splitting Multiple Access (RSMA) has recently emerged as a promising non-orthogonal multiple access technique, offering enhanced flexibility in interference management for next-generation wireless communications [5]. The core concept of RSMA is to handle multi-user interference by partially decoding it and partially treating it as noise [5]. In RSMA, the BS divides the UE signals into common and private components and transmits the combined signal using superposition coding (SC). At the receiver end, the common stream is first decoded while considering all private streams as interference. Once decoded, the common stream is removed from the received signal through successive interference cancellation (SIC). Subsequently, each UE decodes its private stream while treating private streams from other UEs as interference. This adaptive interference management strategy enables RSMA to serve as an intermediate approach between space-division multiple access (SDMA), which entirely treats multi-user interference as noise, and non-orthogonal multiple access (NOMA), which fully decodes interference.
@@ I-A1 FD-BS-based wireless systems
Motivated by the substantial gain brought by the FD communication system, several works have been reported in the literature investigating the interaction between FD and other emerging technologies. For example, the authors in [2] proposed a multi-cell FD system incorporating reconfigurable intelligent surface (RIS) where they showed that the system could significantly improve cell edge users performance. Meanwhile, the authors in [6] presented a real-world evaluation of the FD system with mmWave communication and measured the impact of the SI channel. The paper in [7] explored the optimization of long-term energy efficiency for the FD cell-free massive MIMO system and the formulated problem was solved utilizing a reinforcement learning-based approach. Moving on, the authors in [8] studied the effects of joint transmit and receive antenna selection problems on FD MIMO networks’ performance. In this context, they formulated a sum-rate maximization problem, and the problem was solved using a generalized Bender’s decomposition-based algorithm. On the other hand, the authors in [9] studied a joint antenna and Internet of Things device scheduling problem for FD MIMO wireless-powered communication network over time-varying fading channels. First, they formulated an optimization problem to maximize the average sum rate of Internet of Things devices while satisfying their minimum average data rate requirements, and then they solved the problem by proposing a scheduling algorithm based on Lagrangian duality and the stochastic optimization theory.
@@ I-A2 RSMA/NOMA-assisted FD-BS systems
Inspired by the potential of RSMA/NOMA and FD, several research works have been performed incorporating these two technologies in different scenarios. The authors in [10] investigated an FD-integrated RSMA scheme for improved SE and energy efficiency performance and compared it to the conventional power-domain schemes. In this regard, the authors formulated a multi-objective optimization problem that aimed to jointly maximize energy efficiency and SE by jointly optimizing transmit power budget and minimum rate under the assumption of a channel state information error model. Then, the multi-objective problem was converted into a single-objective optimization problem using the weighted sum method, and successive convex approximation (SCA) and the S-procedure were utilized to solve the formulated problem. Another work presented a model consisting of FD RIS and NOMA with the consideration of hardware impairment [11]. In this regard, they proposed an iterative algorithm that optimized the transmit power at the BS and UL users and the passive beamforming at the RIS.
@@ I-A3 MA-enabled MISO/MIMO/FD systems
MA technology was first introduced in [12], where a thorough study was conducted to highlight its promising potential compared to traditional FPA systems, which suffer from limited flexibility in the spatial domain. Since then, researchers have been exploring innovative system models incorporating various technologies to further amplify its advantages. Multiple-input single-output (MISO)/MIMO technologies, cutting-edge advancements in 5G networks, and key enablers for next-generation wireless systems have proven to increase data throughput and enhance system capacity. As a result, several studies have shifted focus toward MA-enabled MISO/MIMO communication systems. The authors in [13] examined a MIMO system empowered by MA technology at the transmitter’s and receiver’s sides. To characterize the capacity in the model, an optimization problem was formulated to maximize the sum rate by jointly determining the covariance of the transmit signal along with the MAs positions. Besides, in [14], the authors introduced a graph-based technique for optimizing MA positions, contrasting with conventional optimization methods. Their approach involved dividing the transmit area into discrete points, converting the continuous optimization of antenna positions into a problem of selecting the best sampling points based on local channel data. Moving on, the authors in [3] studied an FD-based MA system focusing on the perspective of physical layer security. Particularly, in their model, a BS equipped with MAs operates in FD mode and transmits artificial noise (AN) to simultaneously protect UL and DL transmission. Then, they formulated a sum of secrecy rates maximization problem by jointly optimizing MA positions, the transmit, receive, and AN beamformers at the BS, and the UL powers. Finally, the optimization problem was solved utilizing the AO algorithm by decomposing the main problem into multiple sub-problems. For the optimization of beamforming vectors and UL power, SCA was utilized, and for MA position optimization, a particle swarm-based approach was proposed. Meanwhile, the authors in [15], examined an FD-integrated sensing and communication-based system with MAs to promote communication capability with guaranteed sensing performance via jointly designing beamforming, power allocation, receiving filters, and MA configuration towards maximizing the sum rate. The optimization problem was later solved using the majorization-minimization method.
@@ I-A4 RSMA-based MA-enabled wireless systems
In order to enhance connectivity and SE in upcoming wireless networks, various studies have begun to concentrate on strengthening MA-enabled networks using non-orthogonal transmission methods, particularly NOMA and RSMA. The authors of [16] analyze an RSMA-equipped model in which the BS is equipped with a set of linear MAs. Motivated by the advantages of MA technology in terms of efficient exploitation of spatial diversity, they formulated an optimization problem to maximize the achievable sum rate by jointly determining the beamforming design at the BS, the MA positions and the common stream portion for the different users. Due to the non-convex nature of the problem, a two-stage, coarse-to-fine-grained searching algorithm was adopted to produce effective solutions. In [17], the authors examined a short packet Ultra-Reliable and Low-Latency Communication (URLLC) system that leveraged RSMA as a multiple access technique. To overcome the high cost of deploying a large number of antennas at the BS and to produce more flexible beamforming design, MA technology is deployed at the BS. With the aim of maximizing the sum rate, they presented an optimization problem by jointly optimizing the MA positions along with the MA positions while respecting the URLLC requirements. Then, an effective AO strategy is illustrated to handle the high coupling between the different optimization variables. Utilizing RSMA technology also, the work in [18] investigated a DL coordinated multi-point scenario where a set of BSs are jointly serving a set of users acquiring MA technology. The authors formulated an optimization problem with the objective of maximizing the achievable sum rate by determining the optimal beamforming vectors at the cooperating BSs, the MA positions, and the common stream portions while satisfying the quality of service (QoS) requirements. Due to the non-convex nature of the problem stemming from the high interdependence between the variables, a meta-learning-based algorithm that operates without pre-training was deployed in the solution approach.
@@ I-B Contributions
To the best of our knowledge, no prior study has investigated the integration of FD-BS-RSMA with MA technology. In particular, we leverage MA technology to examine its effectiveness in mitigating SI and intra-cell interference in FD-BS-RSMA system. Particularly, in a traditional FD-BS system with FPA, residual SI and intra-cell interference primarily handled by either designing beamforming and power allocation [2] or by utilizing effective access technique such as RSMA [10]. However, utilizing MA with FD-BS along with RSMA provides another degree of freedom to reduce the residual SI impact and intra-cell interference and thus maximizing sum rate. In this context, we analyze three distinct scenarios that incorporate MAs in an FD-BS-RSMA system. In scenario 1, multiple UEs, each equipped with a single MA, are served by an FD BS equipped with multiple transmit and receive FPAs. In the second scenario, we consider a system consisting of multiple UEs, each equipped with a single FPA, while the FD BS is equipped with multiple transmit and receive MAs. Finally, in the third scenario, we consider both the BS and UEs are equipped with MAs, where the BS is equipped with multiple transmit and receive MAs, and each UE has single MA. For all scenarios, we conduct a comprehensive analysis of MA’s role in suppressing SI and intra-cell interference, as well as its impact on total sum rate, DL sum rate, and UL sum rate. To the best of our knowledge, this is the first study in the literature to provide a detail investigation of MA technology’s effectiveness in managing interference in FD system.
@@
  • The formulated optimization problem is highly non-convex in nature and difficult to tackle using conventional optimization methods due to the high coupling among the different optimization variables. Hence, we propose a gradient-based meta-learning (GML) approach which can handle the optimization problem in a discrete manner by optimizing each variable in a different neural network and having the flexibility of gradient sharing among them. It should be noted that GML has already proven its efficacy in solving large-scale non-convex optimization problems [19].

### v3

@@ Abstract
Full-duplex (FD) radios at base stations (BS)s have gained significant interest because of their ability to simultaneously transmit and receive signals on the same frequency band. However, FD communication is hindered by self-interference (SI) and intra-cell interference caused by simultaneous uplink (UL) transmissions affecting downlink (DL) reception. Movable antenna (MA) technology has recently emerged as a promising solution to mitigate interference by dynamically adjusting antenna positions within the transmitter or receiver region, enhancing multiplexing gain and spectral efficiency (SE). In traditional FD systems with fixed-position antennas (FPAs), residual SI and intra-cell interference are mitigated either through beamforming and power allocation or by leveraging the inherent interference management strategy of multiple access schemes like rate-splitting multiple access (RSMA). Utilizing MA in FD systems adds an extra degree of freedom to effectively mitigate interference in conjunction with other strategies. In this context, we integrate MA with FD systems to evaluate and investigate its effectiveness in handling SI and intra-cell interference. Moreover, we utilize RSMA as our multiple access technique in both UL and DL transmission. To achieve the full potential of the system, we evaluate three different scenarios with FD-BS-RSMA with MAs where our goal is to maximize the total sum rate of the system by jointly optimizing the transmitting and receiving beamforming vectors, UL user equipment (UE) transmission power, MA positions, and common stream split ratio of RSMA while satisfying the minimum data rate requirements of all UEs, common stream rate constraint, power budget requirements of BS and UL UEs, and inter-MA distance. The formulated optimization problem is highly non-convex in nature, and hence, we propose a gradient-based meta-learning (GML) approach, which can handle the non-convexity in a discrete manner by optimizing each variable in a different neural network. Our simulation results show that introducing MA with FD-BS-RSMA system not only improves the overall SE but also effectively tackles SI and intra-cell interference. Our proposed three scenarios, which integrate MA with the FD-BS-RSMA system, achieve performance improvements ranging from 41% to 240% compared to traditional FPA systems.
@@ I Introduction
With the emergence of sixth-generation (6G) mobile communication systems, there is a pressing demand to explore communication technologies that offer high spectral efficiency (SE), reliability, energy efficiency and latency [10258345]. In addition, the continuous surge in data-intensive applications places highly demanding requirements on wireless communication systems. However, due to limited spectrum resources, enhancing the communication rate using traditional half-duplex (HD) systems remains a challenge. In order to fulfill these stringent requirements and challenges, several promising technologies have emerged, such as millimeter wave (mmWave) communication, massive multiple-input multiple output (mMIMO) and full-duplex (FD) communications [10258345]. Among these technologies, FD communication has garnered significant interest because of its superior SE and co-time co-frequency transceiving capability compared to its HD counterpart [10315060, 10158724].
@@
In FD communications, multiple terminals can simultaneously transmit and receive information within the same frequency band [ding2024movable]. In particular, FD communication is capable of doubling up the SE by accommodating both uplink (UL) and downlink (DL) communications at the same time-frequency resource. However, allocating transmit and receive signals to operate in the same frequency and time slot inherently leads to significant interference issues. Specifically, the most critical issue is residual self-interference (SI) from the transmit antennas of the base station (BS) to the receive antennas, which could be much higher than the received signal. In addition, residual SI and DL user equipment (UE) experience significant intra-cell interference caused by UL UE transmissions, which greatly limits the potential gains of using FD technology. Hence, FD technology will greatly suffer quite a loss if the interferences are not handled in an efficient manner.
@@
Recently, movable antenna (MA) has emerged as a contender technology that is capable of effectively reducing interference in the MIMO system by positioning each MA in the transmitter/receiver region to shift from areas of high interference to locations with minimal interference, resulting in higher multiplexing gain and higher SE [zhu2023movable]. Moreover, by moving the antenna positions in a certain region, MA adds a new degree of freedom to exploit the benefit of the spatial domain. Specifically, unlike traditional fixed-position antennas (FPAs), MAs can be dynamically adjusted within a spatial region to optimize channel conditions and enhance communication performance. In particular, an MA is linked to the radio frequency (RF) chain via a flexible cable, enabling it to be repositioned within a spatial region with the help of a driver. Although traditional MIMO systems using antenna selection can achieve spatial degrees of freedom, they generally require a large number of antennas to cover the entire spatial region and fully leverage spatial diversity and multiplexing gains [zhu2023movable]. In contrast, a much smaller number of MAs, even a single MA, can fully exploit these degrees of freedom because of their flexible positioning.
@@
Another crucial aspect of enhancing SE lies in the careful selection of multiple access techniques. In the past couple of years, Rate-Splitting Multiple Access (RSMA) has recently emerged as a promising non-orthogonal multiple access technique, offering enhanced flexibility in interference management for next-generation wireless communications [mao2022rate]. The core concept of RSMA is to handle multi-user interference by partially decoding it and partially treating it as noise [mao2022rate]. In RSMA, the BS divides the UE signals into common and private components and transmits the combined signal using superposition coding (SC). At the receiver end, the common stream is first decoded while considering all private streams as interference. Once decoded, the common stream is removed from the received signal through successive interference cancellation (SIC). Subsequently, each UE decodes its private stream while treating private streams from other UEs as interference. This adaptive interference management strategy enables RSMA to serve as an intermediate approach between space-division multiple access (SDMA), which entirely treats multi-user interference as noise, and non-orthogonal multiple access (NOMA), which fully decodes interference.
@@ I-A1 FD-BS-based wireless systems
Motivated by the substantial gain brought by the FD communication system, several works have been reported in the literature investigating the interaction between FD and other emerging technologies. For example, the authors in [10315060] proposed a multi-cell FD system incorporating reconfigurable intelligent surface (RIS) where they showed that the system could significantly improve cell edge users performance. Meanwhile, the authors in [10477869] presented a real-world evaluation of the FD system with mmWave communication and measured the impact of the SI channel. The paper in [10770280] explored the optimization of long-term energy efficiency for the FD cell-free massive MIMO system and the formulated problem was solved utilizing a reinforcement learning-based approach. Moving on, the authors in [9448430] studied the effects of joint transmit and receive antenna selection problems on FD MIMO networks’ performance. In this context, they formulated a sum-rate maximization problem, and the problem was solved using a generalized Bender’s decomposition-based algorithm. On the other hand, the authors in [park2022joint] studied a joint antenna and Internet of Things device scheduling problem for FD MIMO wireless-powered communication network over time-varying fading channels. First, they formulated an optimization problem to maximize the average sum rate of Internet of Things devices while satisfying their minimum average data rate requirements, and then they solved the problem by proposing a scheduling algorithm based on Lagrangian duality and the stochastic optimization theory.
@@ I-A2 RSMA/NOMA-assisted FD-BS systems
Inspired by the potential of RSMA/NOMA and FD, several research works have been performed incorporating these two technologies in different scenarios. The authors in [allu2024robust] investigated an FD-integrated RSMA scheme for improved SE and energy efficiency performance and compared it to the conventional power-domain schemes. In this regard, the authors formulated a multi-objective optimization problem that aimed to jointly maximize energy efficiency and SE by jointly optimizing transmit power budget and minimum rate under the assumption of a channel state information error model. Then, the multi-objective problem was converted into a single-objective optimization problem using the weighted sum method, and successive convex approximation (SCA) and the S-procedure was utilized to solve the formulated problem. Another work presented a model consisting of FD RIS and NOMA with the consideration of hardware impairment [mead2024hardware]. In this regard, they proposed an iterative algorithm that optimized the transmit power at the BS and UL users and the passive beamforming at the RIS. An RSMA FD-based vehicular-to-everything network has been presented in [10663924], which is also aided by RIS to maximize the SE. The formulated non-convex problem is solved utilizing a deep reinforcement learning-based optimization algorithm. Another work has presented a model consisting of FD and RIS and NOMA with the consideration of hardware impairment [mead2024hardware]. In this regard, they proposed an iterative algorithm that optimizes the transmit power at the BS and UL users and the passive beamforming at the RIS.
@@ I-A3 MA-enabled MISO/MIMO/FD systems
MA technology was first introduced in [10318061], where a thorough study was conducted to highlight its promising potential compared to traditional FPA systems, which suffer from limited flexibility in the spatial domain. Since then, researchers have been exploring innovative system models incorporating various technologies to further amplify its advantages. Multiple-input single-output (MISO)/MIMO technologies, cutting-edge advancements in 5G networks, and key enablers for next-generation wireless systems have proven to increase data throughput and enhance system capacity. As a result, several studies have shifted focus toward MA-enabled MISO/MIMO communication systems. The authors in [10243545] examined a MIMO system empowered by MA technology at the transmitter’s and receiver’s sides. To characterize the capacity in the model, an optimization problem was formulated to maximize the sum rate by jointly determining the covariance of the transmit signal along with the MAs positions. Besides, in [10508218], the authors introduced a graph-based technique for optimizing MA positions, contrasting with conventional optimization methods. Their approach involved dividing the transmit area into discrete points, converting the continuous optimization of antenna positions into a problem of selecting the best sampling points based on local channel data. Moving on, the authors in [ding2024movable] studied an FD-based MA system focusing on the perspective of physical layer security. Particularly, in their model, a BS equipped with MAs operates in FD mode and transmits artificial noise (AN) to simultaneously protect UL and DL transmission. Then, they formulated a sum of secrecy rates maximization problem by jointly optimizing MA positions, the transmit, receive, and AN beamformers at the BS, and the UL powers. Finally, the optimization problem was solved utilizing the alternating optimization (AO) algorithm by decomposing the main problem into multiple sub-problems. For the optimization of beamforming vectors and UL power, SCA was utilized, and for MA position optimization, a particle swarm-based approach was proposed. Meanwhile, the authors in [guo2024movable], examined an FD-integrated sensing and communication-based system with MAs to promote communication capability with guaranteed sensing performance via jointly designing beamforming, power allocation, and receiving filters, and MA configuration towards maximizing the sum rate. The optimization problem was later solved using the majorization-minimization method.
@@ I-A4 RSMA-based MA-enabled wireless systems
In order to enhance connectivity and SE in upcoming wireless networks, various studies have begun to concentrate on strengthening MA-enabled networks using non-orthogonal transmission methods, particularly NOMA and RSMA. The authors of [zhang2024sumratemaximizationmovable] analyzed an RSMA-equipped model in which the BS is equipped with a set of linear MAs. Motivated by the advantages of MA technology in terms of efficient exploitation of spatial diversity, they formulated an optimization problem to maximize the achievable sum rate by jointly determining the beamforming design at the BS, the MA positions, and the common stream portion for the different users. Due to the non-convex nature of the problem, a two-stage, coarse-to-fine-grained searching algorithm was adopted to produce effective solutions. In [10834523], the authors examined a short packet Ultra-Reliable and Low-Latency Communication (URLLC) system that leveraged RSMA as a multiple access technique. To overcome the high cost of deploying a large number of antennas at the BS and to produce a more flexible beamforming design, MA technology is deployed at the BS. With the aim of maximizing the sum rate, they presented an optimization problem by jointly optimizing the MA positions along with the MA positions while respecting the URLLC requirements. Then, an effective AO strategy is illustrated to handle the high coupling between the different optimization variables. Utilizing RSMA technology, the work in [amhaz2025enhancingcomprsmaperformancemovable] investigated a DL coordinated multi-point scenario where a set of BSs are jointly serving a set of users acquiring MA technology. The authors formulated an optimization problem with the objective of maximizing the achievable sum rate by determining the optimal beamforming vectors at the cooperating BSs, the MA positions, and the common stream portions while satisfying the quality of service (QoS) requirements. Due to the non-convex nature of the problem stemming from the high interdependence between the variables, a meta-learning-based algorithm that operates without pre-training was deployed in the solution approach.
@@ I-B Contributions
To the best of our knowledge, no prior study has investigated the integration of FD-BS-RSMA with MA technology. In particular, we leverage MA technology to examine its effectiveness in mitigating SI and intra-cell interference in FD-BS-RSMA system. Particularly, in a traditional FD-BS system with FPA, residual SI and intra-cell interference primarily handled by either designing beamforming and power allocation [10315060] or by utilizing effective access technique such as RSMA [allu2024robust]. However, utilizing MA with FD-BS along with RSMA provides another degree of freedom to reduce the residual SI impact and intra-cell interference and thus maximizing sum rate. In this context, we analyze three distinct scenarios that incorporate MAs in an FD-BS-RSMA system. In scenario 1, multiple UEs, each equipped with a single MA, are served by an FD BS equipped with multiple transmit and receive FPAs. In the second scenario, we consider a system consisting of multiple UEs, each equipped with a single FPA, while the FD BS is equipped with multiple transmit and receive MAs. Finally, in the third scenario, we consider both the BS and UEs are equipped with MAs, where the BS is equipped with multiple transmit and receive MAs, and each UE has single MA. For all scenarios, we conduct a comprehensive analysis of MA’s role in suppressing SI and intra-cell interference, as well as its impact on total sum rate, DL sum rate, and UL sum rate. To the best of our knowledge, this is the first study in the literature to provide a detail investigation of MA technology’s effectiveness in managing interference in FD system.
@@
  • The formulated optimization problem is highly non-convex in nature and difficult to tackle using conventional optimization methods due to the high coupling among the different optimization variables. Hence, we propose a gradient-based meta-learning (GML) approach which can handle the optimization problem in a discrete manner by optimizing each variable in a different neural network and having the flexibility of gradient sharing among them. It should be noted that GML has already proven its efficacy in solving large-scale non-convex optimization problems [zhu2024robust].